Amultimodestep-indexfiberhascoreand cladding refractive indices of 1,475 and 1.458 respectively. If it has a core radius of 30 μm and operates at a wavelength of 1250 nm.



– What should be the core radius of the fiber if it is to operate as a single mode fiber

1
Expert's answer
2022-03-24T16:36:50-0400

Now we know that

N.A="\sqrt{n_1^2-n_2^2}"


"NA=\sqrt{1.475^2-1.458^2}=0.2232"

"\lambda_c=\frac{2\pi a(NA)}{V_{max}}"

"a=\frac{\lambda_c V_{max}}{2\pi (NA)}"

"a=\frac{1250\times10^{-9}\times2.405 }{2\times3.14 \times0.2232}"

"a=2.14\times10^{-6}m"